# A comprehensive survey on deep active learning and its applications in medical image analysis

### v1

@@ Abstract
Deep learning has achieved widespread success in medical image analysis, leading to an increasing demand for large-scale expert-annotated medical image datasets. Yet, the high cost of annotating medical images severely hampers the development of deep learning in this field. To reduce annotation costs, active learning aims to select the most informative samples for annotation and train high-performance models with as few labeled samples as possible. In this survey, we review the core methods of active learning, including the evaluation of informativeness and sampling strategy. For the first time, we provide a detailed summary of the integration of active learning with other label-efficient techniques, such as semi-supervised, self-supervised learning, and so on. Additionally, we also highlight active learning works that are specifically tailored to medical image analysis. In the end, we offer our perspectives on the future trends and challenges of active learning and its applications in medical image analysis.
@@ 1 Introduction
Medical imaging visualizes anatomical structures and pathological processes. It also offers crucial information in lesion detection, diagnosis, treatment planning, and surgical intervention. In recent years, the rise of artificial intelligence (AI) has led to significant success in medical image analysis. The AI-powered systems for medical image analysis have not only approached but even exceeded the performance of human experts in certain clinical tasks. Notable examples include skin cancer classification (Esteva et al., 2017), lung cancer screening with CT scans (Ardila et al., 2019), polyp detection during colonoscopy (Wang et al., 2018), and prostate cancer tissue detection in whole-slide images (Tolkach et al., 2020). Therefore, these AI-powered systems can be integrated into existing clinical workflows, which helps to improve diagnostic accuracy for clinical experts (Sim et al., 2020) and support less-experienced clinicians (Tschandl et al., 2020).
Deep learning (DL) models serve as the core of these AI-powered systems for learning complex patterns from raw images and generalizing them to more unseen cases. The success of DL often relies on large-scale human-annotated datasets. For example, the ImageNet dataset (Deng et al., 2009) contains tens of millions of labeled images, and it’s widely used in developing DL models for computer vision. The size of medical image datasets keeps expanding, but it is still relatively smaller than that of natural image datasets. For example, the brain tumor segmentation dataset BraTS (Menze et al., 2014; Baid et al., 2021) consists of 3D multi-sequence MRI scans. The BraTS dataset expanded from 65 patients in 2013 to over 1,200 in 2021. The latter is equivalent to more than 700,000 annotated 2D images. However, the high annotation cost limits the construction of large-scale medical image datasets, mainly reflected in the following two aspects:
Figure 1: Overall framework of this survey.
@@
2. The high bar for medical image annotation leads to high costs. In computer vision, tasks like object detection and segmentation also require many fine-grained annotations. However, the widespread use of crowdsourcing platforms has significantly reduced the cost of obtaining high-quality annotations in these tasks (Kovashka et al., 2016). However, crowdsourcing platforms have certain limitations in annotating medical images. Firstly, annotating medical images demands both medical knowledge and clinical expertise. Some complex cases even require discussions among multiple senior experts. Secondly, even in some relatively simple tasks, crowdsourcing workers tend to provide annotations of poorer quality than professional annotators in medical image analysis. For example, results in Rädsch et al. (2023) supported the conclusion above in annotating the segmentation mask of surgical instruments. Finally, crowdsourcing platforms may also raise privacy concerns (Rajpurkar et al., 2022). In summary, high-quality annotations often require the involvement of experienced doctors, which inherently increases the annotation cost of medical images.
@@
The high annotation cost is one of the major bottlenecks of DL in medical image analysis. Active learning (AL) is considered one of the most effective solutions for reducing annotation costs. The main idea of AL is to select the most informative samples for annotation and then train a model with these samples in a supervised way. In the general practice of AL, annotating a part of the dataset could reach comparable performance of annotating all samples. As a result, AL saves the annotation costs by querying as few informative samples for annotation as possible. Specifically, we refer to the AL works focusing on training a deep model as deep active learning.
@@
Reviewing AL works in medical image analysis is essential for reducing annotation costs. There are already some surveys on AL in machine learning or computer vision. Settles (2009) provided a general introduction and comprehensive review of AL works in the machine learning era. After the advent of DL, Ren et al. (2021) reviewed the development of deep active learning and its applications in computer vision and natural language processing. Liu et al. (2022b) summarized the model-driven and data-driven sample selectors in deep active learning. Zhan et al. (2022) reimplemented high-impact works in deep active learning with fair comparisons. Takezoe et al. (2023) reviewed recent developments of deep active learning in computer vision and its industrial applications. Regarding related surveys in medical image analysis, Budd et al. (2021) investigated the role of humans in developing and deploying DL in medical image analysis, where AL is considered a vital part of this process. In Tajbakhsh et al. (2020), AL was one of the solutions for training high-performance medical image segmentation models with imperfect annotation. As one of the methods in label-efficient deep learning for medical image analysis, Jin et al. (2023a) summarized AL methods from model and data uncertainty.
@@
However, the surveys mentioned above have certain limitations. First, new ideas and methods are constantly emerging with the rapid development of deep active learning. Thus, a more comprehensive survey of AL is needed to cover the latest advancements. Second, a recent trend is combining AL with other label-efficient techniques, which is also highlighted as a future direction by related surveys (Takezoe et al., 2023; Budd et al., 2021). However, existing surveys still lack summaries and discussions on this topic. Finally, the high annotation cost emphasizes the increased significance of AL in medical image analysis, yet related reviews still lack comprehensiveness in this regard.
Figure 2: Illustration of the process of active learning.
@@
This survey comprehensively reviews AL for medical image analysis, including core methods, integration with other label-efficient techniques, and AL works tailored to medical image analysis. We first searched relevant papers on Google Scholar and arXiv platforms using the keyword “Active Learning” and expanded the search scope through citations. It should be noted that the included papers in this survey mainly belong to the fields of medical image analysis and computer vision. AL works on language, time series, tabular data, and graphs are less emphasized. Additionally, most works in this survey are published in top-tier journals (including TPAMI, TMI, MedIA, JBHI, etc.) and conferences (including CVPR, ICCV, ECCV, ICML, ICLR, NeurIPS, MICCAI, ISBI, MIDL, etc.). As a result, this survey involves nearly 164 relevant AL works with 234 references. The contributions of this paper are summarized as follows:
@@
  • Through an exhaustive literature search, we provide a comprehensive review and a novel taxonomy for AL works, especially those focusing on medical image analysis.
@@
The rest of this survey is organized as follows: §2 introduces problem settings and mathematical formulation of AL, §3 discusses the core methods of AL, including evaluation of informativeness (§3.1 & §3.2) and sampling strategies (§3.3), §4 reviews the integration of AL with other label-efficient techniques, §5 summarizes AL works tailored to medical image analysis. We discuss existing challenges and future directions of AL in §6 and conclude the whole paper in §7. The overall framework of this survey is shown in Fig. 1.
@@ 2 Problem Settings and Formulations of Active Learning
Most AL works follow pool-based active learning, which draw samples from a large pool of unlabeled data and requests oracle (e.g., doctors) for annotations. Moreover, if multiple samples are selected for labeling at once, we can further call this setting “batch-mode”. Deep active learning is in batch-mode by default since retraining the model every time a sample is labeled is impractical. Also, one labeled sample may not necessarily result in significant performance improvement. Therefore, unless otherwise specified, all works in this survey follow the setting of batch-mode pool-based active learning.
@@
1. Sample Selection: In the $t$-th round of annotation, $1\leq t\leq T$, an acquisition function $A$ is used to evaluate the informativeness of each sample in the unlabeled pool $D_{t}^{u}$. Then, a batch of samples is selected with a certain sampling strategy $S$. Specifically, the queried dataset of $t$-th round $D_{t}^{q}$ is constructed as follow:
$D_{t}^{q}=\underset{D_{t}^{q}\subset D_{t}^{u}}{S}\left(\underset{x\in D_{t}^{u}}{A}\left(x,f_{\theta_{t-1}}\right),b\right)$ (1)
@@
2. Annotation by Oracle: After sample selection, the queried set $D_{t}^{q}$ is sent to oracle (e.g., doctors) for annotation, and newly labeled samples are added into the labeled dataset $D_{t}^{l}$. The update of $D_{t}^{l}$ is as follow:
@@
$D_{t}^{u}=D_{t-1}^{u}\backslash\{(x,y)|x\in D_{t}^{q}\}$ (3)
3. DL Model Training: After oracle annotation, we train the deep model using the labeled set of this round $D_{t}^{l}$ in a fully supervised manner. The deep model $f_{\theta_{t}}$ is trained on $D_{t}^{l}$ to obtain the optimal parameters $\theta_{t}$ for round $t$. The mathematical formulation is as follows:
@@
where $\mathcal{L}$ represents the loss function.
4. Repeat steps 1 to 3 until the annotation budget limit is reached.
It is worth noting that the model needs proper initialization to start the AL process. If the initial model $f_{\theta_{0}}$ is randomly initialized, it could only produce meaningless informativeness. To address this issue, most AL works randomly choose some samples as initially labeled dataset $D_{0}^{l}$ and train $f_{\theta_{0}}$ upon $D_{0}^{l}$. For more details on better initialization of AL using pre-trained models, please refer to §4.2.
@@ 3 Core Methods of Active Learning
In this survey, we consider the evaluation of informativeness and sampling strategy as the core methods of AL. Informativeness represents the value of annotating each sample. Higher informativeness indicates a higher priority to request these samples for labeling. Typical metrics of informativeness include uncertainty based on model prediction and representativeness based on data distribution. On the other hand, the sampling strategy is used to select a small number of unlabeled samples for annotation based on their informativeness metrics. Common sampling strategies include top-k selection and clustering, etc. Unlike previous surveys, we explicitly define sampling strategies as core methods of AL for the first time. The rationale is that if a perfect informativeness metric existed, one could simply select the samples with the highest value of the informativeness metric according to the annotation budget. However, current informativeness metrics are more or less flawed to some extent. For instance, redundant and class-imbalanced queries are common issues in AL. We need specific sampling strategies to mitigate the issues arising from imperfect informativeness metrics. In this section, we reviewed uncertainty (§3.1), informativeness (§3.2) and sampling strategy (§3.3). Additionally, we provide a summarization of all works in this survey. Methods and basic metrics of calculating uncertainty or representativeness and sampling strategies are detailed in Table 4.
@@ 3.1 Evaluation of Informativeness: Uncertainty
Uncertainty is used to assess the reliability of model predictions, with higher uncertainty indicating the model may be more prone to errors (Kendall and Gal, 2017). During sample selection in AL, it is challenging to determine the correctness of the model’s predictions. Nevertheless, we can highlight samples where the model is prone to making errors by uncertainty estimation. Uncertain samples often contain knowledge about the model that has not yet been mastered. Annotating them can improve the performance. Therefore, uncertainty has become one of the most commonly used informativeness metrics in AL.
The most straightforward uncertainty metrics are based on prediction probabilities, including least confidence (Lewis and Catlett, 1994), entropy (Joshi et al., 2009), margin (Roth and Small, 2006), and mean variance (Gal et al., 2017). These metrics have been widely used in traditional AL, and their formulations are detailed in Table 1. Confidence is the probability of the highest predicted class. We often adopt the least confidence to measure uncertainty, meaning lower confidence indicates higher uncertainty. Entropy is the most common measure of uncertainty, with higher entropy indicating higher uncertainty. The margin represents the difference between the highest and the second-highest predicted probability, with a larger margin indicating greater uncertainty. Besides, a larger mean variance indicates higher uncertainty.
Table 1: Formulations of Uncertainty Metrics based on Prediction Probability. $x$ stands for sample, $f$ is the deep model, while $C$ is the number of classes.
Name Equations
Prediction Probability $p=\text{Softmax}\left(f\left(x\right)\right)\in\mathbb{R}^{C},p=\left[p_{1},p_{2},\cdots,p_{C}\right]$
@@
Mean Variance $-\frac{1}{C}\sum_{i=1}^{C}\left(p_{i}-\bar{p}\right)^{2},\bar{p}=-\frac{1}{C}\sum_{i=1}^{C}p_{i}$
The above uncertainty metrics only require a single forward pass in deep learning. However, due to the notorious issue of over-confidence in deep neural networks (Guo et al., 2017), they cannot be directly transferred to deep AL. In deep learning, over-confidence refers to the model having excessively high confidence in its predictions, even though they might not be accurate. Over-confidence results in high confidence (e.g., 0.99) of the wrong class for misclassified samples. For uncertain samples, it may lead to extreme confidence (e.g., 0.99 or 0.01) instead of normal one (e.g., 0.6 or 0.4) as it should. Over-confidence may cause distorted uncertainty since it affects the predicted probabilities for all classes. This section divides the uncertainty-based AL into multiple inference, disagreement-based uncertainty, uncertainty-aware models, gradient-based uncertainty, adversarial-based uncertainty, and performance estimation. The first three are primarily based on prediction probabilities, while the latter three mostly employ other statistics of deep models for uncertainty estimation. The taxonomy of uncertainty-based AL is shown in Fig. 3.
@@
3.1.1 Multiple Inferences

### v3

# A comprehensive survey on deep active learning in medical image analysis

@@ Abstract
Deep learning has achieved widespread success in medical image analysis, leading to an increasing demand for large-scale expert-annotated medical image datasets. Yet, the high cost of annotating medical images severely hampers the development of deep learning in this field. To reduce annotation costs, active learning aims to select the most informative samples for annotation and train high-performance models with as few labeled samples as possible. In this survey, we review the core methods of active learning, including the evaluation of informativeness and sampling strategy. For the first time, we provide a detailed summary of the integration of active learning with other label-efficient techniques, such as semi-supervised, self-supervised learning, and so on. We also summarize active learning works that are specifically tailored to medical image analysis. Additionally, we conduct a thorough comparative analysis of the performance of different AL methods in medical image analysis with experiments. In the end, we offer our perspectives on the future trends and challenges of active learning and its applications in medical image analysis.
@@ 1 Introduction
Medical imaging visualizes anatomical structures and pathological processes. It also offers crucial information in lesion detection, diagnosis, treatment planning, and surgical intervention. In recent years, the rise of artificial intelligence (AI) has led to significant success in medical image analysis. The AI-powered systems for medical image analysis have approached the performance of human experts in certain clinical tasks. Notable examples include skin cancer classification (Esteva et al., 2017), lung cancer screening with CT (Ardila et al., 2019), polyp detection during colonoscopy (Wang et al., 2018), and prostate cancer detection in whole-slide images (Tolkach et al., 2020). Therefore, these AI-powered systems can be integrated into existing clinical workflows, which helps to improve diagnostic accuracy for clinical experts (Sim et al., 2020) and support less-experienced clinicians (Tschandl et al., 2020).
Deep learning (DL) models serve as the core of these AI-powered systems for learning complex patterns from raw images and generalizing them to more unseen cases. Leveraging their robust feature extraction and generalization capabilities, DL models have also achieved remarkable success in the field of medical image analysis (Zhou et al., 2021a). The success of DL often relies on large-scale human-annotated datasets. For example, the ImageNet dataset (Deng et al., 2009) contains tens of millions of labeled images, and it’s widely used in developing DL models for computer vision (CV). The size of medical image datasets keeps expanding, but it is still relatively smaller than that of natural image datasets. For example, the brain tumor segmentation dataset BraTS consists of multi-sequence 3D MRI scans. The BraTS dataset expanded from 65 patients in 2013 (Menze et al., 2014) to over 1,200 in 2021 (Baid et al., 2021). The latter is equivalent to more than 700,000 annotated 2D images  [^1] BraTS 2021 dataset has scans with annotations available for 1,251 patients which all have the same spatial shape. Each case features four MRI sequences, and each sequence contains 155 2D axial slices. Thus, it contains a total of $1,251\times 4\times 155=775,620$ slices.. However, the high annotation cost limits the construction of large-scale medical image datasets, mainly reflected in the following two aspects:
Figure 1: Illustration of the process of active learning.
@@
2. The high bar for medical image annotation leads to high costs. In CV, tasks like object detection and segmentation also require fine-grained annotations. However, the widespread use of crowdsourcing platforms has significantly reduced the cost of obtaining high-quality annotations in these tasks (Kovashka et al., 2016). However, crowdsourcing platforms have certain limitations in medical image annotation. Firstly, annotating medical images demands both medical knowledge and clinical expertise. Complex cases even require discussions among multiple senior experts. Secondly, even in relatively simple tasks, crowdsourcing workers tend to provide annotations of poorer quality than professional annotators in medical image analysis. For example, results in Rädsch et al. (2023) supported the conclusion above in annotating the segmentation mask of surgical instruments. Crowdsourcing platforms could also raise privacy concerns (Rajpurkar et al., 2022). Nevertheless, we will face new challenges when the annotators change from crowdsourcing workers to clinical experts. First of all, recruiting doctors for annotation is very expensive. For instance, a radiologist usually takes about 60 minutes to manually segment brain tumors per patient in its multi-sequence MRI volumes(Menze et al., 2014). And the median hourly rate of a radiologist is \$219 in the US [^2] https://www.salary.com/tools/salary-calculator/radiologist-hourly. Besides, to minimize individual bias for certain scenarios, it is common to have a doctor annotate the same case multiple times or have multiple doctors annotate it. Multiple annotation rounds and annotators introduce intra- and inter-annotator variability and handling such variabilities leads to additional annotation costs (Karimi et al., 2020). In summary, high-quality annotations often require the involvement of experienced doctors, which inherently increases the annotation cost of medical images.
@@
The high annotation cost is one of the major bottlenecks of DL in medical image analysis. Active learning (AL) is considered one of the most effective solutions for reducing annotation costs. The main idea of AL is to select the most informative samples for annotation and then train a model with these samples in a supervised way. In the general practice of AL, annotating a part of the dataset could reach comparable performance of annotating all samples. As a result, AL saves the annotation costs by querying as few informative samples for annotation as possible. The process of AL is illustrated in Fig.1, which we will detail in §2. Specifically, we refer to the AL works focusing on training a deep model as deep active learning.
@@
Reviewing AL works in medical image analysis is essential for reducing annotation costs. Budd et al. (2021) investigated the role of humans in developing and deploying DL in medical image analysis, where AL is considered a vital part of this process. In Tajbakhsh et al. (2020), AL was one of the solutions for training high-performance medical image segmentation models with imperfect annotation. As one of the methods in label-efficient deep learning for medical image analysis, Jin et al. (2023a) summarized AL methods from model and data uncertainty. There are also several surveys on AL in machine learning or CV. Settles (2009) provided a general introduction and comprehensive review of AL works in the machine learning era. After the advent of DL, Ren et al. (2021) reviewed the development of deep active learning and its applications in CV and natural language processing. Liu et al. (2022) summarized the model-driven and data-driven sample selectors in deep active learning. Zhan et al. (2022) reimplemented high-impact works in deep active learning with fair comparisons. Takezoe et al. (2023) reviewed recent developments of deep active learning in CV and its industrial applications.
@@
However, the surveys mentioned above have certain limitations. Firstly, new ideas and methods are constantly emerging with the rapid development of deep active learning. Thus, a more comprehensive survey of AL is needed to cover the latest advancements. Secondly, a recent trend is combining AL with other label-efficient techniques, which is also highlighted as a future direction by related surveys (Takezoe et al., 2023; Budd et al., 2021). However, existing surveys still lack summaries and discussions on this topic. Thirdly, limited surveys have evaluated the performance of different AL methods on the medical imaging dataset, indicating a near absence of such efforts. Finally, the high annotation cost emphasizes the increased significance of AL in medical image analysis, yet related reviews still lack comprehensiveness in this regard.
@@
This survey comprehensively reviews AL for medical image analysis, including core methods, integration with other label-efficient techniques, and AL works tailored to medical image analysis. We first searched relevant papers on Google Scholar and arXiv using the keyword “Active Learning” and expanded the search scope through citations. The included papers in this survey mainly belong to the field of medical image analysis. It should be noted that some important works of AL in the general CV field are also included since the development of AL in medical image analysis is influenced by the advance of AL in CV. Ignoring these works would flaw the logic and taxonomy of this survey. To balance the AL works of different fields, we first present the seminal works in each subsection, which may include works in the general CV field, and then provide a detailed review of the AL papers related to medical image analysis within this category. Additionally, most works in this survey are published in top-tier journals (e.g., TPAMI, TMI, MedIA, TBME, JBHI, etc.) and conferences (e.g., CVPR, ICCV, ECCV, ICML, ICLR, NeurIPS, MICCAI, ISBI, MIDL, etc.). As a result, this survey involves nearly 164 relevant AL works with 234 references. The contributions of this paper are summarized as follows:
@@
  • Through an exhaustive literature search, we provide a comprehensive survey and a novel taxonomy for AL works, especially those focusing on medical image analysis.
@@
  • For the sake of promoting research and contributing to the community, this survey evaluated the performance of several popular AL methods on multiple medical imaging datasets. The codes are also made public for better reproducibility.
The rest of this survey is organized as follows: §2 introduces problem settings and mathematical formulation of AL, §3 discusses the core methods of AL, including evaluation of informativeness (§3.1 & §3.2) and sampling strategies (§3.3), §4 reviews the integration of AL with other label-efficient techniques, §5 summarizes AL works tailored to medical image analysis. The experimental settings, results, and analysis are in §6. We discuss existing challenges and future directions of AL in §7 and conclude the whole paper in §8. The overall framework of this survey is shown in Fig. 2.
@@ 2 Problem Settings and Formulations of Active Learning
Most AL works follow pool-based active learning, which draw samples from a large pool of unlabeled data and requests oracle (e.g., doctors) for annotations. Moreover, if multiple samples are selected for labeling at once, we can further call this setting “batch-mode”. Deep active learning is in batch-mode by default since retraining the model every time a sample is labeled is impractical. Also, one labeled sample does not necessarily result in significant performance improvement. Therefore, unless otherwise specified, all works in this survey follow the setting of batch-mode pool-based active learning.
@@
(1) Sample Selection: In the $t$-th round of annotation, $1\leq t\leq T$, an informativeness function $I$ is used to evaluate the informativeness of each sample in the unlabeled pool $D_{t}^{u}$. Then, a batch of samples is selected with a certain sampling strategy $S$. In medical image analysis, active learning selects a batch of images most of the time (i.e., image-wise selection). In this survey, the sampling unit of AL selection is an image (could be 2D or 3D) unless specifically stated. However, with the development of AL, region-wise (§4.3, §4.5.2) or slice-wise annotations (§5.2.1) are adopted in AL. Please refer to these sections for more details.
Specifically, the queried dataset of $t$-th round $D_{t}^{q}$ is constructed as follow:
$D_{t}^{q}=\underset{D_{t}^{q}\subset D_{t}^{u}}{S}\left(\underset{x\in D_{t}^{u}}{I}\left(x,f_{\theta_{t-1}}\right),b\right)$ (1)
@@
(2) Annotation by Oracle: After sample selection, the queried set $D_{t}^{q}$ is sent to oracle (e.g., doctors) for annotation, and newly labeled samples are added into the labeled dataset $D_{t}^{l}$. The update of $D_{t}^{l}$ is as follow:
@@
$D_{t}^{u}=D_{t-1}^{u}\backslash\{x|x\in D_{t}^{q}\}$ (3)
It is worth noting that some current works combine active learning with interactive segmentation. In interactive segmentation, the model assists experts in annotation, thereby reducing the difficulty of the annotation process. For more details, please refer to section §5.2.4.
(3) DL Model Training: After oracle annotation, we train the deep model using the labeled set of this round $D_{t}^{l}$ in a fully supervised manner. The deep model $f_{\theta_{t}}$ is trained on $D_{t}^{l}$ to obtain the optimal parameters $\theta_{t}$ for round $t$. The mathematical formulation is as follows:
@@
where $\mathcal{L}(f_{\theta}(x),y)$ represents the loss function and it could be rewritten as $\mathcal{L}(x,y;\theta)$ for simplicity.
(4) Repeat steps 1 to 3 until the annotation budget limit or the expected performance is reached. Recently, some works adopted the one-shot fashion in active learning, which performed sample selection without multiple rounds. Please refer to §5.2.2.
It is worth noting that the model needs proper initialization to start the AL process. If the initial model $f_{\theta_{0}}$ is randomly initialized, it could only produce meaningless informativeness. To address this issue, most AL works randomly choose a set of samples as initially labeled dataset $D_{0}^{l}$ and train $f_{\theta_{0}}$ upon $D_{0}^{l}$. For more details on better initialization of AL using pre-trained models, please refer to §4.2.
@@ 3 Core Methods of Active Learning
In this survey, we consider the evaluation of informativeness and sampling strategy as the core methods of AL. Informativeness represents the value of annotating each sample. Higher informativeness often indicates a higher priority to request these samples for labeling. Typical metrics of informativeness include uncertainty and representativeness. Based on the informativeness scores, a certain sampling strategy is used to select a small number of unlabeled samples for annotation. Most AL works simply ranked these samples by their informativeness metrics and selected the highest ones according to the annotation budget (i.e., top-k selection). However, current informativeness scores are more or less flawed and they may cause issues like redundancy or class imbalance among the queried samples. Therefore, we need more advanced sampling strategies to mitigate these issues arising from imperfect informativeness metrics.
In this section, we reviewed two major informativeness metrics, including uncertainty (§3.1) and representativeness (§3.2), and sampling strategy (§3.3). As a unique contribution of this survey, we, for the first time, explicitly define sampling strategies as core methods of AL and review how to design a better sampling strategy in AL. Additionally, we provide a summarization of all the cited AL works in this survey. Methods and basic metrics of uncertainty or representativeness and sampling strategies are detailed in Table 4.
@@ 3.1 Evaluation of Informativeness: Uncertainty
Table 1: Formulations of uncertainty metrics based on prediction probability in active learning. In the equations column, $x$ stands for sample, $f$ is the deep model, while $C$ is the number of classes. In the direction column, $\uparrow$ means higher values indicating higher uncertainty, while $\downarrow$ means lower values indicating higher uncertainty.
Names Equations Direction
Prediction Probability $\mathbf{p}=\text{Softmax}\left(f_{\theta}\left(x\right)\right)\in\mathbb{R}^{C},$ $\mathbf{p}=\left[p_{1},p_{2},\cdots,p_{C}\right]$
@@
Mean Variance (Gal et al., 2017) $-\frac{1}{C}\sum_{i=1}^{C}\left(p_{i}-\bar{p}\right)^{2},\bar{p}=-\frac{1}{C}\sum_{i=1}^{C}p_{i}$uparrow
Despite great progress has been made in medical image analysis, safety and interpretability are still unsolved problems for deploying DL models in real-world clinical practice. Due to the high variability of medical images and the limited training data, the predictions of DL models are not reliable and trusted. Correctly assessing and quantifying uncertainty in medical image analysis would allow models to alert ambiguities, artifacts, and unseen patterns in the data (Ghesu et al., 2021; Linmans et al., 2023). Such properties of uncertainty are helpful in AL since novel patterns in the unlabeled samples can be identified by the uncertainty. Therefore, uncertainty is frequently used in active learning as an informative metric. In the AL query, samples with higher uncertainty are considered hard and more likely to be misclassified by the current model. Annotating and training on these samples helps the model learn new patterns and improve performance.
Tracing back to the cause of uncertain predictions, the uncertainty can be mainly separated into two types: aleatoric uncertainty (AU) and epistemic uncertainty (EU) (Kendall and Gal, 2017). AU (i.e., data uncertainty) captures the noisy observations in data, such as motion artifacts of MRI or metal artifacts of CT in medical image analysis. AU cannot be reduced by acquiring more data. A high EU (i.e., model uncertainty) indicates that the samples contain knowledge that has not yet been mastered by the model. Therefore, the EU can be reduced by involving more data. However, most of the AL works did not consider the separation of AU and EU. So, the terminology of uncertainty in AL mainly refers to prediction uncertainty, which is the composition of AU and EU. This is because explicitly separating AU and EU is usually very difficult and may not bring much benefit in the practice of AL (Kahl et al., 2024). In this survey, unless explicitly stated otherwise, all uncertainties refer to predictive uncertainties, meaning that no separation between AU and EU has been performed.
The most straightforward uncertainty metrics in deep AL are based on prediction probabilities with a single forward pass. These metrics have been widely used in AL since the machine learning era, and their formulations are detailed in Table 1. However, directly transferring them to deep AL would be challenging due to the notorious issue of over-confidence in deep neural networks (Mehrtash et al., 2020; Guo et al., 2017). Over-confidence refers to the model having excessively high confidence in its predictions, even though they might not be accurate. It could result in high confidence (e.g., 0.99) of the wrong class for misclassified samples. For uncertain samples, it leads to extreme confidence (e.g., 0.99 or 0.01) instead of normal one (e.g., 0.6 or 0.4) as it should. As a result, over-confidence distorts uncertainty estimation since it affects the predicted probabilities for all classes.
This section divides the uncertainty-based AL into multiple inference, gradient-based uncertainty, performance estimation, uncertainty-aware models, and adversarial-based uncertainty. The taxonomy of uncertainty-based AL is shown in Fig. 3.
@@
3.1.1 Uncertainty with Multiple Inferences